OpenStudy (anonymous):

i need help with the problem 20-3h=2 The answer is supposed to be 6 but I cant figure out how to get there. I understand to solve it you do the order of operations backwards but no matter what order i try I can not get the correct answer. could someone please help me understand? Thanks. i will medal and fan!

OpenStudy (twistrose):

is it suppose to be positive or negative

OpenStudy (anonymous):

positive

OpenStudy (twistrose):

20-3h=2 -20 -20 -3h=-18 /-3 /-3 6 Does this help?

OpenStudy (anonymous):

but isn't the 20 supposed to be added?

OpenStudy (twistrose):

no its suppose to be subtracted because its positive (+)

OpenStudy (anonymous):

oh, okay so when you 'split' the problem up the 3h is negative because of the subtraction sight right before it?

OpenStudy (twistrose):

correct

OpenStudy (anonymous):

Thanks a ton!!

OpenStudy (twistrose):

your welcome :)